An agency surveyed people worldwide, asking them, Do you live in a household with more than one car? Of the 1399 respondentPlease help asap!!
Answer #1

Here, , n1 = 1399 , n2 = 809
p1cap = 0.2194 , p2cap = 0.7219


Standard Error, sigma(p1cap - p2cap),
SE = sqrt(p1cap * (1-p1cap)/n1 + p2cap * (1-p2cap)/n2)
SE = sqrt(0.2194 * (1-0.2194)/1399 + 0.7219*(1-0.7219)/809)
SE = 0.0193

For 0.99 CI, z-value = 2.58
Confidence Interval,
CI = (p1cap - p2cap - z*SE, p1cap - p2cap + z*SE)
CI = (0.2194 - 0.7219 - 2.58*0.0193, 0.2194 - 0.7219 + 2.58*0.0193)
CI = (-0.552 , -0.453)

Lower limit = -0.552